Shift Summary - Day - 20 june 1999

Douglas A. Jensen (djensen@fnal.gov)
Sun, 20 Jun 1999 22:04:09 -0500 (CDT)

Shift Summary for day shift, 20/jun/99

On Shift: Tony Barker, Doug Jensen

And, Ed, Hogan, Rick K, ...

Hours of beam delivered at all intensities:
Hours of beam delivered at nominal intensity:
Hours of E832 data (special runs not included):
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:changing tapes, accelerator down

========================================================================
Runs ended this shift:
Run 13668 - 23.9MM events - BEAM_832 - 36.4K vacuum 2pi0

Tapes brought to FCC this shift:, none
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Took over running experiment.
About 10:45, see beam is off by -16mm, soon moved to -19 mm.
NM2EU problem. Shut down for repairs. Beam back about 12:43

CsI channel 449 was off - shifting up to 5%, rms 2.5%. Turns out
that as Ed discovered, doing a pipeline reset solved the problem.
Was true at start of shift, probably was true for esentially all
of run 13668.

The LN2 dewar is empty. Trd Folks interested ?

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13668 07:00 to 10:45, good run we thought. But need to study
sdaq sem means to know


13669 aborted
13670 11:30 to ... open stop at 12:55,beam tuning continues.


========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Douglas Jensen DJensen@fnal.gov x8194